# Towards Universal Soccer Video Understanding

### v2

@@ 4.4 Downstream Tasks
After the abovementioned training, MatchVision can now serve as a versatile visual encoder, to map the soccer video segments into visual features ($\mathcal{F}_{\mathcal{V}}$), for training task-specific heads $\Psi=\{\Psi_{\mathrm{cls}},\Psi_{\mathrm{Cmt}},\Psi_{\mathrm{Foul}}\}$ for different downstream tasks, including: (i) event classification, (ii) commentary generation, and (iii) foul recognition.
Event Classification. Similar to the supervised training mentioned above, we concatenate a learnable [cls] token to aggregate the frame-wise visual features through temporal self-attention. This token is then fed into a linear classifier for event classification. The event classification head ($\Psi_{\mathrm{cls}}$) can be trained with a cross-entropy loss with the visual encoder frozen.
Commentary Generation. On commentary generation, we follow the design in MatchTime [40]. Concretely, the commentary generation head ($\Psi_{\mathrm{Cmt}}$) employs a Perceiver [24] aggregator to consolidate visual features, which are then projected by a trainable MLP, serving as prefix embeddings for a large language model (LLM). Subsequently, an off-the-shelf LLM decodes these embeddings into textual commentary. We adopt the negative log-likelihood loss, commonly used for auto-regressive next-token prediction.
Foul Recognition. As outlined in [21], the foul recognition task takes multi-view videos from the same scene as inputs, with each sample annotated with a foul class (8 types) and severity (4 levels). We encode these multi-view videos with MatchVision, and then aggregate the extracted features into a single feature vector, via either max or average pooling, following the common practice. Subsequently, the foul recognition head ($\Psi_{\mathrm{Foul}}$) employs a shared MLP and two task-specific linear classifiers, to predict foul type and severity, respectively. Similar to event classification, we use the combination of cross-entropy losses on the foul type and severity classification to jointly train $\Psi_{\mathrm{Foul}}$.
@@
Table 2: Quantitative Results on Event Classification and Commentary Generation. Here, SN, MT, and SR represent curated SoccerNet-v2 [9], MatchTime [40], and SoccerReplay-1988, respectively. Moreover, B, M, R-L, and C refer to BLEU, METEOR, ROUGE-L, and CIDEr metrics, respectively. Within each unit, we denote the best performance in RED and the second-best performance in BLUE.
@@ 5 Experiments
This section begins with the implementation details in Section 5.1; followed by quantitative evaluations across multiple downstream tasks in Section 5.2; then, we conduct ablation experiments on our established SoccerReplay-test benchmark to investigate the effectiveness of the proposed dataset and model in Section 5.3; finally, we provide qualitative results for comparison in Section 5.4.
@@ 5.1 Implementation Details
In our experiments, all video segments are sampled at 1FPS centered around the annotated timestamps, capturing a 30-second window for each sample. Each frame is resized to $224\times 224$ pixels to serve as inputs. We initialize the embedding layer, spatial attention layers, aggregation layer, and text encoder of our MatchVision with the pretrained weights from SigLIP Base-16 [57] and adopt LLaMA-3 (8B) [11] as the off-the-shelf LLM decoder for commentary generation head. All experiments are conducted on 4×4$\times$4 × Nvidia H800 GPUs, using AdamW [33] as the optimizer. Next, we will elaborate on the training and evaluation details about visual encoder pretraining and downstream tasks.
@@
Downstream Tasks. In all downstream tasks, unless otherwise specified, we use the frozen visual encoder for feature extraction and only train the task-specific heads with a learning rate of $1\times 10^{-4}$ for 30 epochs. The batch size for event classification, commentary generation, and foul recognition is set to 40, 32, and 8, respectively. We adopt specific evaluation metrics for these three tasks: (i) For event classification, we use the top-1/3/5 classification accuracy; (ii) For commentary generation, we employ several commonly-used language evaluation metrics, including BLEU [35], METEOR [2], ROUGE-L [32], and CIDEr [48]. (iii) For foul recognition, we follow the common practice, and report top-1/2 and top-1 accuracy for the foul type and severity classification, respectively.
Benchmarks & Baselines. To ensure fair and reliable comparisons with existing work, we evaluate event classification (24 types) on 100 matches from curated SoccerNet-v2 [9] test set, commentary generation on 49 matches from SN-Caption-test-align benchmark manually aligned in [40]; and foul recognition evaluation on MVFoul [21]. We consider various baselines: for the first two tasks, this includes off-the-shelf visual encoders such as ResNet [20], C3D [45], I3D [4], CLIP [38], SigLIP [57], and InternVideo [49], along with Baidu [59] and SigLIP [57] trained under various settings. For foul recognition, we follow previous work and adopt ResNet [20], R(2+1)D [46], and MViT [12] jointly finetuned with classifiers, as baselines.
@@
Table 3: Quantitative Results on Multi-view Fouls Recognition. Our frozen MatchVision encoder can achieve comparable performance with other jointly finetuned visual encoders.
@@ 5.2 Quantitative Evaluation
As presented in Table 2, we can draw two observations on event classification and commentary generation: (i) visual encoders trained on soccer data substantially outperform the off-the-shelf general encoders, ResNet, C3D, I3D, CLIP, and InternVideo, underscoring the necessity of building specific foundation models for soccer understanding; (ii) almost all visual encoders, across all training settings, significantly benefit from SoccerReplay-1988, emphasizing the value of constructing high-quality, large-scale data for soccer video understanding. Next, we will delve deeper into each task to discuss the results.
Event Classification. With the same training strategies and data, MatchVision considerably outperforms other methods in classification accuracy, demonstrating the superiority of our architecture, which effectively leverages spatiotemporal features within soccer videos. Moreover, models trained via supervised training excel others, primarily because the pre-training task shares the same objectives as the downstream event classification task.
Commentary Generation. Visual encoders trained with visual-language contrastive learning exhibit better commentary generation performance than those trained with supervised classification, as this strategy more effectively learns the correlations between visual and textual features. Additionally, the MatchVision encoder trained solely on SoccerNet slightly underperforms the Baidu [59] encoder, but incorporating our SoccerReplay-1988 dataset enables it to outperform on most metrics. This indicates that MatchVision can take the benefit of large-scale datasets. Finally, a hybrid training approach, starting with supervised classification followed by visual-language contrastive learning, enables MatchVision to achieve optimal performance. This indicates that learning coarse-grained tasks such as classification provides a foundation for fine-grained tasks like commentary generation, and that fully leveraging data can unleash the potential of soccer understanding.
Foul Recognition. As demonstrated in Table 3, MatchVision achieves performance comparable to jointly finetuned state-of-the-art methods in foul recognition, even with a frozen visual encoder. This highlights that MatchVision can effectively learn substantial knowledge from large-scale soccer data and adapt seamlessly to downstream tasks.
@@
Table 4: Ablation Studies on Event Classification. We explore the impact of various training settings of our MatchVision encoder on the SoccerReplay-test benchmark. Here, Sup., Contra., and SR refer to supervised classification, visual-language contrastive learning, and the SoccerReplay-1988 dataset, respectively.
Figure 4: Qualitative Results for Event Classification and Commentary Generation. Here, “w/o SR” and “w/ SR” indicate models trained without and with the SoccerReplay-1988 dataset, respectively. Incorporating the SoccerReplay-1988 dataset improves event classification accuracy. For commentary generation, this enriched training data enables the commentary generation head to demonstrate notable advantages in several aspects: (a) more detailed descriptions, (b) greater linguistic variety, (c) higher accuracy in event depiction, (d) better adherence to updated rules, and (e) improved specificity in scenario response.
@@ 5.3 Ablation Studies
We perform ablation experiments on event classification and commentary generation using our SoccerReplay-test benchmark. These experiments validate the effectiveness of our proposed dataset and model, while establishing a baseline for future evaluations on this benchmark.
Event Classification. We evaluate event classification on 300 matches from our SoccerReplay-test benchmark using the MatchVision visual encoder pretrained with various strategies. Features are extracted by MatchVision and processed with a learnable aggregation layer and a linear classifier. The default training set is our curated SoccerNet-pro. As shown in Table 4, integrating our proposed SoccerReplay-1988 for training results in significant performance improvements across all pretraining strategies, yielding the significance of our dataset. Additionally, supervised classification outperforms visual-language contrastive learning and hybrid pretraining. This is due to its closer alignment with downstream event classification, and the scale of event annotations is far larger than that of textual commentaries, further confirming the significant benefits of data scaling for boosting soccer understanding.
Commentary Generation. With the pretrained MatchVision encoder, we train the commentary generation head on the MatchTime [40] and SoccerReplay-1988 datasets using various training strategies. By default, only the Perceiver [24] aggregation layer and projection layer within the head are trained. For joint training with the LLM decoder, considering computational costs, we incorporate LoRA [23] layers while keeping the original LLM layers frozen. As shown in Table 5, incorporating SoccerReplay-1988 significantly improves performance on all metrics, confirming the substantial advantages of our proposed dataset. This performance gap also reflects the challenges of our established benchmark, which features diverse vocabulary, richer semantics, and updated soccer rules. Additionally, jointly finetuning the visual encoder and the LLM decoder provides a feasible approach for further improvements.
@@ 5.4 Qualitative Comparisons
As depicted in Figure 4, we present the qualitative results of MatchVision on the new SoccerReplay-test benchmark, comparing models pretrained without and with the SoccerReplay-1988 dataset. For event classification, incorporating our new dataset improves accuracy, and even in the mis-classified cases, the results remain contextually relevant. For commentary generation, after hybrid training on SoccerReplay-1988, MatchVision is capable of producing richer, more detailed textual commentary, demonstrating a deeper understanding of soccer dynamics. More qualitative results are available in Section D.2 of the Appendix.
Table 5: Ablation Studies of Commentary Generation. We investigate the impact of different training strategies and datasets on MatchVision using the SoccerReplay-test benchmark. ‘V’ and ‘L’ denote the visual encoder and the LLM decoder, respectively.
@@ 6 Conclusion
In this paper, we establish a unified, scalable multi-modal framework for soccer understanding. Specifically, we introduce SoccerReplay-1988, the largest and most comprehensive soccer video dataset to date, annotated by an automated curation pipeline. This provides a solid foundation for developing multi-modal soccer understanding models and serves as a more challenging benchmark. Built upon this, we have developed the first visual-language foundation model for soccer, termed MatchVision, which effectively leverages the spatiotemporal information within soccer videos and can be applied to various tasks such as event classification and commentary generation. Extensive experiments demonstrate the superiority of our proposed model, with MatchVision achieving state-of-the-art performance on both existing benchmarks and our newly established one. We believe this work will set a viable, universal paradigm for future research in sports understanding.
@@ Appendix A SoccerReplay-1988 Dataset Details
In this section, we provide additional details of our SoccerReplay-1988 dataset. Specifically, Section A.1 elaborates on the structure and format of the dataset; Section A.2 presents statistical information and analyses of the dataset; and Section A.3 describes the methodology to generate the event labels within the dataset.
@@ A.1 Dataset Format
The SoccerReplay-1988 dataset consists of match videos, descriptions of events, and related match information of 1988 soccer matches. Each match includes two mkv video files (1-half and 2-half), covering the match from the initial kick-off to the final whistle. Additionally, a json file is accompanied by encapsulating detailed information, including event descriptions and comprehensive match background structured as follows:
Match Information provides background details of the match, including the competing teams, final results, and match contexts, such as the start time, team formations, and venue details, as illustrated below:
@@
    "Image URL": "https://static.flashsc...",                          # URL to player image
@@ A.2 Additional Dataset Statistics
England Premier League 552
@@
To provide a comprehensive analysis of the SoccerReplay-1988 dataset, we present statistical summaries and visualizations in the following tables and figures. Specifically, Table 6 and 7 illustrate the distribution of the 1988 matches across different leagues and seasons. Figure 5 (a) compares the SoccerReplay-1988 dataset with other soccer datasets [14, 9, 40, 36, 25, 56, 43], highlighting its unique characteristics. Figure 5 (b) depicts the distribution of event labels for the 24 newly defined categories. Furthermore, Figures 5 (c), (d), (e), and (f) present detailed analyses of the commentary data, including frequency distributions, timestamps, and word counts.
@@
A.3 Event Summarization Prompt
Our dataset expands the original 17 event categories to 24 types. For SoccerReplay-1988 and MatchTime [40], LLaMA-3 (70B) [11] is used to analyze commentaries and generate event labels. All commentaries are processed with the following prompt with comprehensive definitions and rich examples to determine their types:

### v3

@@ 4.4 Downstream Tasks
After the pretraining mentioned above, MatchVision can now serve as a versatile visual encoder, to map the soccer video segments into visual features ($\mathcal{F}_{\mathcal{V}}$), for training task-specific heads $\Psi=\{\Psi_{\mathrm{cls}},\Psi_{\mathrm{Cmt}},\Psi_{\mathrm{Foul}}\}$ across different downstream tasks, including: (i) event classification, (ii) commentary generation, and (iii) foul recognition.
Event Classification. Similar to supervised classification above, we concatenate a learnable [cls] token to aggregate frame-wise visual features via temporal self-attention. This token is then fed into a linear classifier for event classification. The event classification head ($\Psi_{\mathrm{cls}}$) is trained with a cross-entropy loss while freezing the visual encoder.
Commentary Generation. We follow the paradigm in MatchTime [41] to generate anonymized textual commentary for soccer video clips. Concretely, the commentary generation head ($\Psi_{\mathrm{Cmt}}$) employs a Perceiver [25] aggregator to consolidate visual features, which are then projected by a trainable MLP, serving as prefix embeddings for a large language model (LLM). Subsequently, an off-the-shelf LLM decodes these embeddings into textual commentary. We adopt the negative log-likelihood loss, commonly used for auto-regressive next-token prediction.
Foul Recognition. As outlined in [22], the foul recognition task takes multi-view videos from the same scene as inputs, with each sample annotated with a foul class (8 types) and severity (4 levels). We encode these multi-view videos with MatchVision, and aggregate the extracted features into a single feature vector, via either max or average pooling, following the common practice. Subsequently, the foul recognition head ($\Psi_{\mathrm{Foul}}$) employs a shared MLP and two task-specific linear classifiers, to predict foul type and severity, respectively. Similar to event classification, we use the combination of cross-entropy losses on the foul type and severity classification to jointly train $\Psi_{\mathrm{Foul}}$.
@@
Table 2: Quantitative Results on Event Classification and Commentary Generation. Here, SN, MT, and SR represent finetuning with curated SoccerNet-v2 [10], MatchTime [41], and SoccerReplay-1988, respectively. B, M, R-L, and C refer to BLEU, METEOR, ROUGE-L, and CIDEr metrics, respectively. Within each unit, we denote the best performance in RED and the second-best performance in BLUE.
@@ 5 Experiments
This section begins with implementation details in Sec. 5.1; followed by quantitative evaluations across downstream tasks in Sec. 5.2; then, we conduct ablation studies on our SoccerReplay-test benchmark to analyze the effectiveness of the proposed dataset and model in Sec. 5.3; finally, we provide qualitative results for comparison in Sec. 5.4.
@@ 5.1 Implementation Details
In our experiments, video segments are sampled at 1FPS around annotated timestamps, capturing a 30-second window for each sample. Frames are resized to $224\times 224$ pixels as inputs. We initialize the embedding layer, spatial attention layers, aggregation layer, and text encoder of MatchVision with pretrained weights from SigLIP Base-16 [60] and adopt LLaMA-3 (8B) [12] as the off-the-shelf LLM decoder for commentary generation. All experiments are conducted on 4×4$\times$4 × Nvidia H800 GPUs with the AdamW [34] optimizer. Next, we elaborate on the training and evaluation details about visual encoder pretraining and downstream tasks.
@@
Downstream Tasks. In all downstream tasks, unless otherwise specified, we use the frozen visual encoder for feature extraction and only train the task-specific heads with a learning rate of $1\times 10^{-4}$ for 30 epochs. The batch sizes for event classification, commentary generation, and foul recognition are set to 40, 32, and 8, respectively. We adopt specific evaluation metrics for these three tasks: (i) For event classification, we use the top-1/3/5 classification accuracy; (ii) For commentary generation, we employ several commonly-used language evaluation metrics, including BLEU [36], METEOR [2], ROUGE-L [33], and CIDEr [49]; (iii) For foul recognition, we follow the common practice, and report top-1/2 and top-1 accuracy for the foul type and severity classification, respectively.
Benchmarks & Baselines. To ensure fair and reliable comparisons with existing work, we evaluate event classification (24 types) on 100 matches from curated SoccerNet-v2 [10] test set; commentary generation on 49 matches from SN-Caption-test-align benchmark manually aligned in [41]; and foul recognition on MVFoul [22]. We consider various baselines: for the first two tasks, this includes off-the-shelf general visual encoders such as ResNet [21], C3D [47], I3D [4], CLIP [39], SigLIP [60], and InternVideo [50], along with Baidu [63] and SigLIP finetuned with soccer-specific data. For foul recognition, we follow previous work [22, 23] and adopt ResNet [21], R(2+1)D [48], and MViT [13] jointly finetuned with classifiers, as baselines.
@@
Table 3: Quantitative Results on Multi-view Foul Recognition. Our frozen MatchVision encoder can achieve comparable performance with other jointly finetuned visual encoders.
@@ 5.2 Quantitative Evaluation
As depicted in Table 2, we draw two observations on event classification and commentary generation: (i) visual encoders trained on soccer data substantially outperform off-the-shelf general encoders (ResNet, C3D, I3D, CLIP, and InternVideo), underscoring the necessity of building specialized models for soccer understanding; (ii) almost all visual encoders, across all training settings, benefit from SoccerReplay-1988, emphasizing the value of constructing large-scale, high-quality data for soccer understanding. Next, we will delve into each task to discuss the results.
Event Classification. With identical training strategies and data, MatchVision considerably outperforms other methods in classification accuracy, demonstrating the superiority of its architecture, which effectively leverages spatiotemporal features within soccer videos. For example, MatchVision achieves a Top-1 accuracy of 82.5%, significantly surpassing SigLIP’s 55.9% under the same training conditions. Moreover, models trained via supervised classification excel others, primarily because the pre-training task shares the same objectives as the downstream event classification task.
Commentary Generation. Visual encoders trained with visual-language contrastive learning exhibit better commentary generation performance than those trained with supervised classification, as this strategy better captures correlations between visual and textual features. Additionally, while MatchVision trained solely on SoccerNet slightly underperforms Baidu [63], incorporating SoccerReplay-1988 enables it to outperform on most metrics. This demonstrates that MatchVision can take advantage of large-scale datasets. Finally, a hybrid training approach, starting with supervised classification followed by visual-language contrastive learning, enables MatchVision to achieve optimal performance. This indicates that learning coarse-grained tasks such as classification provides a foundation for fine-grained tasks like commentary generation, and fully leveraging data unlocks the potential of soccer understanding.
Foul Recognition. As demonstrated in Table 3, MatchVision achieves performance comparable to jointly finetuned state-of-the-art methods in foul recognition, even with a frozen visual encoder. This highlights that MatchVision effectively learns substantial knowledge from large-scale soccer data and adapts seamlessly to downstream tasks. Comparisons with additional baselines from the SoccerNet foul recognition challenges [9] are provided in the Appendix.
@@
Table 4: Ablations on Event Classification. We explore the impact of various training settings of our MatchVision encoder on the SoccerReplay-test benchmark. Here, Sup., Contra., and SR refer to supervised classification, visual-language contrastive learning, and the SoccerReplay-1988 dataset, respectively.
Figure 4: Qualitative Results for Event Classification and Commentary Generation. Here, “w/o SR” and “w/ SR” indicate models trained without and with the SoccerReplay-1988 dataset, respectively. Incorporating SoccerReplay-1988 improves event classification accuracy. Moreover, this enriched training data enables the model to demonstrate several advantages in commentary generation: (a) more detailed descriptions, (b) greater linguistic variety, (c) higher event depiction accuracy, (d) better adherence to updated rules, and (e) improved specificity in scenario response.
@@ 5.3 Ablation Studies
We conduct ablation experiments on event classification and commentary generation using our SoccerReplay-test benchmark. These experiments validate the effectiveness of our proposed dataset and model, while establishing a baseline for future evaluations on this benchmark.
Event Classification. We evaluate event classification on 300 matches from our SoccerReplay-test benchmark using the MatchVision visual encoder pretrained with various strategies. Features are extracted by MatchVision and processed with a learnable aggregation layer and a linear classifier. The default training set is our curated SoccerNet-pro. As shown in Table 4, integrating SoccerReplay-1988 for training results in significant performance improvements across all pretraining strategies, yielding the significance of our dataset. Additionally, supervised classification outperforms visual-language contrastive learning and hybrid pretraining. This is due to its closer alignment with downstream event classification task, and the scale of event annotations is far larger than that of textual commentaries, further confirming the substantial benefits of data scaling for boosting soccer understanding.
Commentary Generation. With the pretrained MatchVision encoder, we train the commentary generation head on the MatchTime [41] and SoccerReplay-1988 datasets using various training strategies. By default, only the Perceiver [25] aggregation layer and projection layer within the head are trained. For joint training with the LLM decoder, considering computational costs, we incorporate LoRA [24] layers while freezing the original LLM layers. As shown in Table 5, incorporating SoccerReplay-1988 significantly improves performance on all metrics, confirming substantial advantages of our proposed dataset. This performance gap also reflects the challenges of our established benchmark, which features diverse vocabulary, richer semantics, and updated soccer rules. Additionally, jointly finetuning the visual encoder and the LLM decoder provides a feasible approach for further improvements.
Table 5: Ablations on Commentary Generation. We investigate the impact of different training strategies and datasets on MatchVision using the SoccerReplay-test benchmark. ‘V’ and ‘L’ denote the visual encoder and the LLM decoder, respectively.
@@ 5.4 Qualitative Comparisons
As depicted in Figure 4, we present qualitative results of MatchVision on the SoccerReplay-test benchmark, comparing models pretrained with and without SoccerReplay-1988. For event classification, incorporating our data improves accuracy, and even in misclassified cases, the results remain contextually relevant. For commentary generation, hybrid training on SoccerReplay-1988 enables MatchVision to produce richer, more detailed textual commentary, reflecting a deeper understanding of soccer dynamics. More qualitative results are available in the Appendix.
@@ 6 Conclusion
In this paper, we establish a unified, scalable multi-modal framework for soccer understanding. Specifically, we introduce SoccerReplay-1988, the largest and most comprehensive soccer video dataset to date, annotated by an automated curation pipeline. This provides a solid foundation for developing soccer understanding models and serves as a more challenging benchmark. Built upon this, we develop MatchVision, an advanced soccer-specific visual encoder, which effectively leverages spatiotemporal information within soccer videos and can be applied to various tasks such as event classification and commentary generation. Extensive experiments demonstrate the superiority of our model, with MatchVision achieving state-of-the-art performance on both existing benchmarks and our newly established one. We believe this work will set a viable, universal paradigm for future research in sports understanding.
Acknowledgments
This work is funded by National Key R&D Program of China (No.2022ZD0161400).
@@ Appendix A SoccerReplay-1988 Dataset Details
In this section, we provide additional details of our SoccerReplay-1988 dataset. Specifically, Sec. A.1 elaborates on the structure and format of the dataset; Sec. A.2 presents statistical information and analyses of the dataset; and Sec. A.3 describes the methodology to automatically generate event labels within the dataset.
@@ A.1 Dataset Format
The SoccerReplay-1988 dataset consists of match videos, descriptions of events, and related game information of 1988 soccer matches. Each match includes two mkv video files (1-half and 2-half), covering the match from the initial kick-off to the final whistle. Additionally, a json file is accompanied by encapsulating detailed information, including event descriptions and comprehensive match backgrounds, structured as follows:
Match Information provides background details of the match, including competing teams, final results, and match contexts, such as start time, team formations, and venue details, as illustrated below:
@@ A.2 Additional Dataset Statistics
England Premier League 552 z
@@
To provide a comprehensive analysis of SoccerReplay-1988, we present statistics and visualizations in the following tables and figures. Concretely, Table 6 and 7 illustrate the distribution of the 1988 matches across different leagues and seasons. Figure 5 (a) compares SoccerReplay-1988 with other soccer datasets [15, 10, 41, 37, 26, 59, 45], highlighting its unique scale. Figure 5 (b) depicts the distribution of event labels for the 24 newly defined categories. Finally, Figures 5 (c), (d), (e), and (f) present detailed analyses of commentary data, including frequency distributions, timestamps, and word counts.
@@
A.3 Event Summarization
Our dataset expands the original 17 event categories in SoccerNet [10] to 24 types. For MatchTime [41] and SoccerReplay-1988, LLaMA-3 (70B) [12] is employed to analyze commentaries and generate corresponding event labels. This is guided by a carefully refined prompt, iteratively improved through manual checks and enriched with comprehensive definitions and examples. Notably, text commentaries unrelated to visual content (e.g., ‘possession ratio is 55:45’) are categorized as ‘statistics and summary’, and excluded from model training and testing. Finally, a random sample of 2% of the data yields 98% manual verification accuracy, confirming the high quality of automated labeling. The entire prompt is presented below: